The area of the earth surface is 530 million km. Approximately 7/10 of the earth surface is water. What fraction is land and How many square km of the earth surface is covered with water.
Answer by Cromlix(4381)

Hi there,
7/10 of 530 million km
= 371 million km (Water)
530 - 371 = 159 million km
159/530 = 3/10 (Land)
Hope this helps :-)